A ball is thrown upward at an angle of 30o from the horizontal. What are the values of the horizontal and vertical components of its velocity?


1
Expert's answer
2021-10-24T18:18:15-0400

Unfortunately, from the condition of the question we don't know the initial velocity of the ball. Let's suppose that the initial velocity of the ball is 10 m/s. Then, we can find  the values of the horizontal and vertical components of its velocity:


"v_{0x}=v_0cos\theta=10\ \dfrac{m}{s}\times cos30^{\circ}=8.66\ \dfrac{m}{s},""v_{0y}=v_0sin\theta=10\ \dfrac{m}{s}\times sin30^{\circ}=5.0\ \dfrac{m}{s}."